If nearly 10 5 coulombs liberate 1 gm-equivalent of aluminium, then the amount of aluminium (equivalent weight 9), deposited through electrolysis in 20 minutes by a current of 50 amp. will be (a) 0.6 gm (b) 0.09 gm (c) 5.4 gm (d) 10.8 gm



If nearly 105 coulombs liberate 1 gm-equivalent of aluminium, then the amount of aluminium (equivalent weight 9), deposited through electrolysis in 20 minutes by a current of 50 amp. will be


        (a)    0.6 gm


        (b)    0.09 gm


        (c)    5.4 gm     


        (d)    10.8 gm

Grade:11

1 Answers

Hrishant Goswami
10 years ago

(c)

m = Zit, 9 = Z × 105, Z = 9 × 10-5

Again,

m = Zit = 9 × 10-5 × 50 × 20 × 60 = 5.4 gm